Parth (parthkohli):

\[\int {3y \over (5y^2 + 4)^2} dy\]

Parth (parthkohli):

\[u = 5y^2 + 4 \]

Parth (parthkohli):

\[du = 10ydy \]

Parth (parthkohli):

\[ydy = {1 \over 10}du \]

Parth (parthkohli):

So then substituting:\[ \int{3y \over u^2}dy\]Next, take 3 out:\[3\int{y \over u^2} dy \]Substituting for \(ydy\) next:\[3\int u^{-2} {1 \over 10}du \]

Parth (parthkohli):

Taking 1/10 out:\[{3 \over 10}\int u^{-2} du \]

Parth (parthkohli):

\[{3 \over 10} \times {u^{-1} \over -1}+c \]?

OpenStudy (anonymous):

Well done

Parth (parthkohli):

Thanks, and then:\[{-3 \over 10} u^{-1} + c \]

Parth (parthkohli):

Substituting back for \(u\).\[{-3 \over 10}\left(5y^2 + 4 \right)^{-1} + c \]

Parth (parthkohli):

Does that become\[{-3 \over 10\left(5y^2 + 4 \right)^{}} \]?

Parth (parthkohli):

+ c

OpenStudy (anonymous):

yes

Parth (parthkohli):

Thank you guys for taking time just for me!

OpenStudy (anonymous):

anytime parth